OpenStudy (anonymous):

how would you find the value of sin pi/12 using a half angle formula?

OpenStudy (ash2326):

eli do you know what's \(sin^2x\)=?

OpenStudy (anonymous):

2sin(x)cos(x)

OpenStudy (ash2326):

no sin^2x= 1-cos^2x

OpenStudy (anonymous):

would the answer be 1/2 srq of 2- srq of 3?

OpenStudy (ash2326):

now let me guide you we have to find \(sin\pi/12\) let's find \(cos \pi/12\) then find \(sin\pi/12\) using = \(\sqrt{1-cos^2 \pi/12}\) we know \[cos 2x= 2cos^2 x-1\] here x = pi/12 so \[cos 2\pi/12= 2cos^2 \pi/12-1\] cos 2pi/12= cos pi/6 eli what's cos pi/6

OpenStudy (phi):

the half-angle formula for sin is \[sin(a/2)= \sqrt{\frac{1}{2}(1-cos(a))} \]

OpenStudy (ash2326):

cos pi/6= \(\frac{\sqrt 3}{2}\) so we get \[ \frac{\sqrt 3}{2}=2 cos^2 \pi/12-1\] so \[ cos^2 \pi/12= \frac{1}{2}(\frac{\sqrt 3}{2}+1)\] now \[sin^2 \pi/12= 1- cos^2 \pi/12\] so we get \[sin^2 \pi/12= 1- \frac{1}{2}(\frac{\sqrt 3}{2}+1)\] simplifying we get \[sin^2 \pi/12= \frac{1}{2} - \frac{\sqrt 3}{4}\] so we get \[sin \frac{\pi}{12}= \sqrt{\frac{1}{2} - \frac{\sqrt 3}{4}}\]